OpenStudy (anonymous):

Two forces act on a point object 100 N and 100 N and the angle between them is 170. Solve for the resultant vector if one of the forces is in the positive x direction.

OpenStudy (anonymous):

How do I solve for this?

OpenStudy (amistre64):

the same way as the last one ... the process is the same, only the value of the vectors change

OpenStudy (anonymous):

@amistre64 is 17.43 the answer?

OpenStudy (souvik):

\[F _{net}=\sqrt{F ^{2}_{1}+F ^{2}_{2}+2F _{1}F _{2}\cos \theta }\]

OpenStudy (amistre64):

its asking for a vector, not a force

OpenStudy (amistre64):

resultant vector = vector 1 + vector 2

OpenStudy (amistre64):

v1 = 100 cos(0) i + 100 sin(0) j v2 = 100 cos(170) i + 100 sin(170) j

OpenStudy (anonymous):

@amistre64 v1=100 v2= -81.12

OpenStudy (anonymous):

resultant vector= 18.88 ?

OpenStudy (amistre64):

that is not what the vectors mean; i and j denote seperate components of a vector

OpenStudy (amistre64):

the vector (3,4) does not equal 7

OpenStudy (amistre64):

the vector (3,4) can be written as: 3i + 4j, which is not 7

OpenStudy (amistre64):

100cos(0) i + 100 sin(0) j + 100 cos(170) i + 100 sin(170) j consider the i and j parts as "variables" and combine "like terms"

OpenStudy (amistre64):

stacking the vectors tends to organize them better, but writting them all out in a line is fine as well

OpenStudy (anonymous):

ok thanks!